a ball is dropped on the ground from a height 10cm .if coifficient of restitution is 0.4, then find the height to which ball is rebound ?
 

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Final velocity would be,
v = root( 0 + 2g*10/100) = > root(2) m/s.
From coeff. of restitution,
v’ = ev = > 0.4*v
Now let h be the height it rebounced then,
0^2 = (0.4v)^2 – 2gh
h = 0.16v^2/2g = > 0.16v^2/20 = > 0.32/20 = > 0.016 meter.